Circular No. 7722

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001em IN UGC 11794 M. Papenkova and W. D. Li, University of California at Berkeley, re-

port the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 18.6) in unfiltered images taken with the Katzman Automatic Imaging Telescope (KAIT) on Sept. 20.3 and 21.3 UT. SN 2001em is located at α = 21^{h}42^{m}23^{s}.66, δ = +12^{o}29′50′′.9 (equinox 2000.0), which is 11′′.1 east and 3′′.5 south of the nucleus of UGC 11794. An unfiltered image taken with the 0.8-m Tenagra II telescope by J. Wray, C. W. Chleborad, and M. Schwartz on Sept. 15.3 already showed a hint of the new object. KAIT images taken on Sept. 5.3 (limiting mag ∼ 19.5) and 10.3 (limiting mag ∼ 18.5) showed nothing at this positon.

COMET 19P/BORRELLY D. Schleicher, Lowell Observatory, writes: “I obtained narrowband

photometry of comet 19P on Sept. 18 and 19 using the Hall 1.1-m telescope at Lowell Observatory, yielding the following averaged production-rate results: log Q(OH) = 28.34; equivalent log Q(water; vectorial) = 28.41; log Afρ = 2.50 (cf. IAUC 7342). The radial fall-off of the dust is significantly steeper than the canonical 1/ρ, with Afρ decreasing by 1.9 times between ρ = 20 000 and 110 000 km.”

COMET C/2000 WM_{1} (LINEAR) Schleicher adds that his narrowband photometry of comet C/2000

WM_{1}, obtained as above on Sept. 18 and 19 (at r = 2.33), yields the following averaged results: log Q(OH) = 27.87; equivalent log Q(water; vectorial) = 27.87; log Afρ = 2.42. No temporal or aperture variations were observed.

COMET P/2001 Q6 (NEAT) Additional astrometry and the following orbital elements, published on

MPEC 2001-S06, indicate this comet (cf. IAUC 7698) to be of short period:

T = 2001 Nov. 9.4373 TT  ω = 43.^{o}3153
e = 0.825605  Ω = 22.1467 }2000.0
q = 1.408422 AU  i = 56.8755
a = 8.076029 AU n^{o} = 0.0429445 P = 22.951 years
2001 September 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green